From a cruise ship travelling due north at 16.5 km/h, a wrecked ship x and an observation tower y are observed in line due east. One hour later the wrecked ship and the tower have bearings S34*40'E and S65*10'E.
Find the distance between the wrecked ship and the tower.
1
Expert's answer
2020-11-29T19:18:49-0500

Explanations & Calculations





  • Refer to the sketch attached
  • "\small x" is the distance to the ship wreck from where the cruise ship was an hour ago & "\small y" is the distance between wrecked ship and the light house.
  • "\small l" is the distance the ship travelled during an hour &

"\qquad\qquad
\begin{aligned}
\small l &= \small ut =16.5kmh^{-1}\times 1h=16.5km\\
\\
\small \theta &= \small 34^040^{'}=34^0+\frac{40}{60}^0=34.667^0\\
\small \alpha &= \small 60^010^{'}=65^0+\frac{10}{60}^0=65.167^0
\end{aligned}"

  • Therefore, by trigonometry

"\qquad\qquad
\begin{aligned}
\small x+y &= \small l\times \tan\alpha=35.655km\\
\small x &= \small l\times \tan\theta=\small 11.411km\\
\small \bold{\therefore y} &= \small \bold{23.244km}
\end{aligned}"